f(x) = 5x and g(x) = x-4
Find: f[g(2)]

Found 2 solutions by stanbon, jim_thompson5910:
Answer by stanbon(75887)

f(x) = 5x and g(x) = x-4
Find:
f[g(2)]= f[2-4] = f(-2) = 5(-2) = -10
====================
Cheers,
Stan H.
-----

Answer by jim_thompson5910(35256)

g(x) = x-4

g(2) = 2-4

g(2) = -2

f(x) = 5x

f(g(2)) = 5*g(2)

f(g(2)) = 5*(-2)

f(g(2)) = -10